the 3 rd term and 7 th term of an A.P are 13 & 33 respy..find the nth term of the ap

the 3rd term and 7th term of an A.P are 13 & 33 respy..find the nth term of the ap
 

Grade:10

3 Answers

viswanath
9 years ago
a+2d=13 and a+6d=33. On solving, a=3,d=5. Tn=a+(n-1)d=5n-2.
CHANDRA KIRAN
9 years ago
Given,
3rd term is 13.So, a+2d=13
7th term is 33.So,a+6d=33
On solving,
a=3,d=5
Tn=a+(n-1)d
Tn=3+5n-5=5n-2
So, the nth term of this A.P is 5n-2.
 
CHANDRA KIRAN
9 years ago
Given,
3rd term is 13.So, a+2d=13
7th term is 33.So,a+6d=33
On solving,
a=3,d=5
Tn=a+(n-1)d
Tn=3+5n-5=5n-2
So, the nth term of this A.P is 5n-2.